OpenStudy (anonymous):

help! can u solve: 9+3tanӨ = 3 tan(sq) Ө +1

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

this is a quadratic equation: arrange them first in higher order like y=Ax^2 + Bx + C 9+3tanӨ = 3 tan(sq) Ө +1 0=3 tan(sq) Ө -3tanӨ -9+1 0=3 tan(sq) Ө -3tanӨ -8, now use the quadratic formula a=3 b= -3 c= -8 -b + sqrt(b^2 - 4 ac X= -------------------- 2a -(-3) + srt( 3^2 - 4(3)(-8)) tan D = ----------------------------- 2(3) tan D+ = 2.2078251 calculator and D+ = 1.1455 radians or 65.63 degrees -(-3) - srt( 3^2 - 4(3)(-8)) tan D = ----------------------------- 2(3) tan D-=- 1.207825128 calc D = arc tan(-1.207825128) = -0,87925 radian =-50.377 degree